OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 1)2 – 16? Identify the x–intercepts by graphing the quadratic equation.

OpenStudy (anonymous):

plug x = 0, u get y intercept plug y = 0, u get x intercept

OpenStudy (anonymous):

y = (x – 1)2 – 16 y = (0 – 1)2 – 16 y = (– 1)2 – 16 y = 1 – 16 = -15 (0,-15) y = (x – 1)2 – 16 0 = x2 - 2x +1 - 16 0 = x2 - 2x -15 0=(x - 5)(x + 3) (5,0) (-3,0)